# MEMIT-Merge: Addressing MEMIT’s Key-Value Conflicts in Same-Subject Batch Editing for LLMs

### v1

@@ Abstract
Zilu Dong, Xiangqing Shen, Rui Xia School of Computer Science and Engineering Nanjing University of Science and Technology, China {zldong, xiangqing.shen, rxia}@njust.edu.cn
@@ 1 Introduction
Knowledge editing techniques aim to update models’ internal knowledge without retraining. As large language models (LLMs) continue to scale up, the prohibitive cost of full model retraining has made knowledge editing increasingly crucial in this domain. Among prevalent editing algorithms, a class of algorithms, termed as “Locate and Edit” methods by Zhang et al. (2024), operates on the fundamental assumption that specific knowledge representations are localized within particular regions of the model’s parameter space, enabling targeted modifications through precise manipulation of these identified regions. MEMIT Meng et al. (2023b), one of the most prominent algorithms in this class, has gained significant attention Li et al. (2024); Fang et al. (2024); Gupta et al. (2024). MEMIT inherits the core architectural feature of ROME Meng et al. (2023a), which localizes knowledge to specific layers and modifies the output linear layer of MLP modules to update knowledge. The distinctive advancement of MEMIT lies in its capability to perform batch-wise mass knowledge editing, enabling simultaneous modification of multiple knowledge instances within a single batch.
@@
However, our investigation reveals a critical limitation in MEMIT: When handling batches with multiple edits that share the same subject (such as “John Smith now plays basketball.” and “John Smith comes from England.” where both statements have the subject “John Smith”), the method will exhibit significant performance degradation. In contrast, edits with distinct subjects (such as “Jack Johnson now plays basketball” and “Paul Morand comes from China” with subjects ‘Jack Johnson” and “Paul Morand”, respectively) maintain stable efficacy. To systematically demonstrate this performance degradation, we constructed two contrastive datasets comprising batches with identical subjects versus fully unique subjects, named distinct-subject and same-subject, respectively. The experimental results are in Fig. 1, where the vertical axis represents efficacy (which means the editing success rate) and the horizontal axis indicates the batch size per edit. The results reveal that MEMIT maintains a high success rate as batch size increases when editing distinct-subject cases, but exhibits significant performance degradation for same-subject cases. More detailed experimental settings can be found in Sec. 5.3. Batch editing of same-subject knowledge is essential in practical applications. For instance, updating multiple attributes about a specific individual (such as occupation, nationality, and affiliations) inherently requires concurrent editing of same-subject facts.
Our analysis reveals that the root cause lies in MEMIT’s key-value modeling framework: when handling edits involving the same subject, identical keys (derived from shared subject representations) are forced to map to divergent values (corresponding to distinct knowledge updates). This results in inconsistencies during batch editing. Specifically, MEMIT treats each knowledge instance as a key-value pair, where the input vector to the MLP’s output linear layer serves as the key, and the desired output becomes the value.  [^1] Note that the key-value here refers to the hidden state and output within the MLP module as described by Meng et al. (2023a), rather than the query, key and value in the attention module. The algorithm modifies the MLP layer by computing target values based on new knowledge objects and updating weights to map existing keys to these new values. However, MEMIT’s per-fact value computation creates inherent conflicts: within a single batch, multiple entries sharing identical keys (from same-subject cases) require different value mappings. As MLP output linear layers cannot produce multiple outputs for identical inputs, these conflicting updates lead to the performance degradation. Furthermore, we observe that the closer the keys are to each other, the more significant the performance degradation.
@@ 2 Related Work
Knowledge editing techniques for large language models (LLMs) primarily fall into two paradigms: non-parametric approaches that preserve original parameters and parametric methods that directly modify model weights. Parametric approaches, while effective for targeted updates, often introduce uncontrolled parameter perturbations that adversely affect unrelated knowledge — a challenge addressed through various constraint mechanisms. The parametric category features two dominant subclasses: One is “Meta-Learning Based Methods”, such as MEND Mitchell et al. (2022) and MALMEN Tan et al. (2024) which train meta-networks using carefully designed datasets containing both unrelated knowledge samples and paraphrased sentences, aiming to enhance generalization while minimizing collateral damage. Another is Locate-and-Edit Methods, which includes techniques such as Knowledge Neuron (KN) Dai et al. (2022), identify critical knowledge storage locations before executing precise edits. ROME Meng et al. (2023a) extends this by incorporating knowledge preservation terms in its optimization objective to maintain model integrity.
Our work builds upon MEMIT Meng et al. (2023b), a state-of-the-art locate-and-edit approach that enables batch knowledge editing through MLP layer modifications. Building on MEMIT, many recent methods have made modifications to parameter update methods during editing or to the architecture and location of the edits. PMET Li et al. (2024) incorporates the output of the attention layer in the calculation of parameter updates. AlphaEdit Fang et al. (2024) improves upon MEMIT’s parameter matrix update method by projecting the update matrix into the null space of the original knowledge to mitigate interference with unrelated knowledge. UNKE Deng et al. (2024) extends structured knowledge editing to unstructured editing.
@@ 3.1 Preliminaries
MEMIT employs a triple $(s,r,o)$ to represent the subject, relation and object of a piece of knowledge to be edited, and constructs a complete sentence based on these elements for the editing process. During the editing process of MEMIT, the key is determined by the subject of the sentence and the part preceding the subject. The value, on the other hand, needs to be a value that enables the model to output the object. Therefore, it is optimized backward from the object of the sentence:
@@ 4.1 Cause Analysis
$AKD^{(l)}=\frac{1}{\binom{|B|}{2}}\sum_{\begin{subarray}{c}e_{1}\in B\\ e_{2}\in B\end{subarray}}||k^{(l)}_{e_{1}}-k^{(l)}_{e_{2}}||_{2}$ (3)
@@
We compute the $AKD$ for all layers of the model at the subject’s last token position. As the degree of subject variation increases across sentences, the $AKD$ value proportionally rises. Conversely, when all sentences share identical subjects, the $AKD$ value remains constant at 0.
We construct sentence batches using predefined templates, where batches sharing the same template exhibited similar $AKD$ values, while distinct templates yielded significantly different $AKD$ measurements. The specific templates and corresponding $AKD$ values are detailed in App. B. For experimental validation, we select three $AKD$ groups (0, 10, 25) and conduct editing tests using Qwen2.5-1.5B-Instruct. As shown in Fig. 3, where $AKD$ values are computed using keys from MEMIT’s final editing layer, the results demonstrate an inverse relationship: lower $AKD$ values correspond to reduced editing success rates. This pattern remains consistent across other $AKD$ values, establishing a statistically significant negative correlation between $AKD$ and editing efficacy.
@@ 5.1 Dataset
As mentioned in the Introduction, we constructed two counterfactual knowledge editing datasets based on Wikidata, one with the same subject and the other with distinct subjects. One dataset, named the “same-subject” dataset, contains 100 editing knowledge triples, all with the subject John Smith. The other dataset, named “distinct-subject”, has knowledge triples with subjects unique to each other, while the relations and objects remain consistent with those in the “same-subject” dataset. For specific details on the construction of the datasets, refer to App. A.
In terms of evaluation metrics, we refer to the metrics used by Meng et al. (2023b), namely Efficacy, Paraphrase, and Specificity. Efficacy is the edit success rate, which measures the probability that the edited model produces the correct answer to the sentences used for editing. Paraphrase is measured in the same way as Efficacy, but it uses paraphrased versions of the sentences used for editing. Specificity measures the probability that facts unrelated to the edit remain consistent before and after the edit.
Although there are currently no off-the-shelf datasets for this purpose, hence the datasets for our experiments were self-constructed, as we previously stated, updating multiple attributes about a specific individual is a highly realistic demand. In addition, changing certain information about an entity sometimes requires a chain of changes to other information about that entity. Therefore, addressing such same-subject issues is necessary in real-world scenarios.
@@ Appendix A Details of Constructing Same-Subject and Distinct-Subject Data
Subsequently, following the dataset metrics in Meng et al. (2023a), we add two types of questions: specificity and paraphrase. For paraphrase questions, we use the same knowledge triples as the edit sentences, but with a different template format: “The name of the relation of subject is object.”. For specificity, there are two types of questions. One is completely unrelated knowledge, for which we use the prompt “The capital city of America is”. The other type has the same relation as the edited knowledge but a different subject. For example, if the edited knowledge is (John, father, Bob), a specificity question could be (Paul, father, Eugène).
@@ Appendix C More Results with Varying Batch Sizes
It shows clearly that MEMIT-based methods suffers from same subject issue, while methods like FT doesn’t.

### v2

@@ Abstract
Zilu Dong  and Xiangqing Shen  and Rui Xia School of Computer Science and Engineering Nanjing University of Science and Technology, China {zldong, xiangqing.shen, rxia}@njust.edu.cn
@@ 1 Introduction
Knowledge editing techniques aim to update models’ internal knowledge without retraining. As large language models (LLMs) continue to scale up, the prohibitive cost of full model retraining has made knowledge editing increasingly crucial in this domain. Among prevalent editing algorithms, a class of algorithms, termed “Locate and Edit” methods by Zhang et al. (2024), operates on the fundamental assumption that specific knowledge representations are localized within particular regions of the model’s parameter space, enabling targeted modifications through precise manipulation of these identified regions. MEMIT Meng et al. (2023), one of the most prominent algorithms in this class, has gained significant attention Li et al. (2024); Fang et al. (2024); Gupta et al. (2024). MEMIT inherits the core architectural feature of ROME Meng et al. (2022), which localizes knowledge to specific layers and modifies the output linear layer of MLP modules to update knowledge. The distinctive advancement of MEMIT lies in its capability to perform batch-wise mass knowledge editing, enabling simultaneous modification of multiple knowledge instances within a single batch.
@@
However, our investigation reveals a critical limitation in MEMIT: When handling batches with multiple edits that share the same subject (such as “John Smith now plays basketball.” and “John Smith comes from England.” where both statements have the subject “John Smith”), the method will exhibit significant performance degradation. In contrast, edits with different subjects (such as “Jack Johnson now plays basketball” and “Paul Morand comes from China” with subjects ‘Jack Johnson” and “Paul Morand”, respectively) maintain stable efficacy. To systematically demonstrate this performance degradation, we constructed two contrastive datasets comprising batches with identical subjects versus fully unique subjects, named distinct-subject and same-subject, respectively. The experimental results are in Fig. 1, where the vertical axis represents efficacy (which means the editing success rate) and the horizontal axis indicates the batch size per edit. The results reveal that MEMIT maintains a high success rate as batch size increases when editing distinct-subject cases, but exhibits significant performance degradation for same-subject cases. More detailed experimental settings can be found in Sec. 5.3. Batch editing of same-subject knowledge is essential in practical applications. For instance, updating multiple attributes about a specific individual (such as occupation, nationality, and affiliations) inherently requires concurrent editing of same-subject facts.
Our analysis reveals that the root cause lies in MEMIT’s key-value modeling framework: when handling edits involving the same subject, identical keys (derived from shared subject representations) are forced to map to divergent values (corresponding to distinct knowledge updates). This results in inconsistencies during batch editing. Specifically, MEMIT treats each knowledge instance as a key-value pair, where the input vector to the MLP’s output linear layer serves as the key, and the desired output becomes the value.  [^1] Note that the key-value here refers to the hidden state and output within the MLP module as described by Meng et al. (2022), rather than the query, key and value in the attention module. The algorithm modifies the MLP layer by computing target values based on new knowledge objects and updating weights to map existing keys to these new values. However, MEMIT’s per-fact value computation creates inherent conflicts: within a single batch, multiple entries sharing identical keys (from same-subject cases) require different value mappings. As MLP output linear layers cannot produce multiple outputs for identical inputs, these conflicting updates lead to the performance degradation. Furthermore, we observe that the closer the keys are to each other, the more significant the performance degradation.
@@ 2 Related Work
Knowledge editing techniques for large language models (LLMs) primarily fall into two paradigms: non-parametric approaches that preserve original parameters and parametric methods that directly modify model weights. Parametric approaches, while effective for targeted updates, often introduce uncontrolled parameter perturbations that adversely affect unrelated knowledge — a challenge addressed through various constraint mechanisms. The parametric category features two dominant subclasses: One is “Meta-Learning Based Methods”, such as MEND Mitchell et al. (2022) and MALMEN Tan et al. (2024) which train meta-networks using carefully designed datasets containing both unrelated knowledge samples and paraphrased sentences, aiming to enhance generalization while minimizing collateral damage. Another is Locate-and-Edit Methods, which includes techniques such as Knowledge Neuron (KN) Dai et al. (2022), identify critical knowledge storage locations before executing precise edits. ROME Meng et al. (2022) extends this by incorporating knowledge preservation terms in its optimization objective to maintain model integrity.
Our work builds upon MEMIT, a state-of-the-art locate-and-edit approach that enables batch knowledge editing through MLP layer modifications. Building on MEMIT, many recent methods have made modifications to parameter update methods during editing or to the architecture and location of the edits. PMET Li et al. (2024) incorporates the output of the attention layer in the calculation of parameter updates. AlphaEdit Fang et al. (2024) improves upon MEMIT’s parameter matrix update method by projecting the update matrix into the null space of the original knowledge to mitigate interference with unrelated knowledge. UNKE Deng et al. (2024) extends structured knowledge editing to unstructured editing.
@@ 3.1 Preliminaries
MEMIT employs a triple $(s,r,o)$ to represent the subject, relation, and object of a piece of knowledge to be edited, and constructs a complete sentence based on these elements for the editing process. During the editing process of MEMIT, the key is determined by the subject of the sentence and the part preceding the subject. The value, on the other hand, needs to be a value that enables the model to output the object. Therefore, it is optimized backward from the object of the sentence:
@@ 4.1 Cause Analysis
$AKD^{(l)}=\frac{1}{\binom{|B|}{2}}\sum_{\begin{subarray}{c}e_{1}\in B\\ e_{2}\in B\end{subarray}}||k^{(l)}_{e_{1}}-k^{(l)}_{e_{2}}||_{2}$
@@
We compute $AKD$ for all layers of the model at the subject’s last token position. As the degree of subject variation increases across sentences, the $AKD$ value proportionally rises. Conversely, when all sentences share identical subjects, the $AKD$ value remains constant at 0.
We construct sentence batches using predefined templates, where batches sharing the same template exhibited similar $AKD$ values, while distinct templates yielded significantly different $AKD$ measurements. The specific templates and the corresponding values of $AKD$ are detailed in App. B. For experimental validation, we select three $AKD$ groups (0, 10, 25) and conduct editing tests using Qwen2.5-1.5B-Instruct. As shown in Fig. 3, where $AKD$ values are computed using keys from MEMIT’s final editing layer, the results demonstrate an inverse relationship: lower $AKD$ values correspond to reduced editing success rates. This pattern remains consistent across other $AKD$ values, establishing a statistically significant negative correlation between $AKD$ and editing efficacy.
@@ 5.1 Dataset
As mentioned in the Introduction section, we constructed two counterfactual knowledge editing datasets based on Wikidata, one with the same subject and the other with distinct subjects. The dataset named “same-subject” dataset, contains 100 editing knowledge triples, all with the subject John Smith. The other dataset, named “distinct-subject”, has knowledge triples with subjects unique to each other, while the relations and objects remain consistent with those in the “same-subject” dataset. For specific details on the construction of the datasets, refer to App. A.
In terms of evaluation metrics, we refer to the metrics used by Meng et al. (2023), namely Efficacy, Paraphrase, and Specificity. Efficacy is the edit success rate, which measures the probability that the edited model produces the correct answer to the sentences used for editing. Paraphrase is measured in the same way as Efficacy, but it uses paraphrased versions of the sentences used for editing. Specificity measures the probability that facts unrelated to the edit remain consistent before and after the edit.
Although we constructed our own datasets due to the lack of specialized data on the same subject in the past, addressing same-subject issue has a practical necessity. Editing multiple attributes of an entity (e.g, updating a person’s profile) is a highly realistic demand. Furthermore, changing certain information about an entity sometimes requires a chain of changes to other information about that entity. All of these make the same subject scenarios essential for practical applications.
@@ Appendix A Details of Constructing Same-Subject and Distinct-Subject Data
Subsequently, following the dataset metrics in Meng et al. (2022), we add two types of questions: specificity and paraphrase. For paraphrase questions, we use the same knowledge triples as the edit sentences, but with a different template format: “The name of the relation of subject is object.”. For specificity, there are two types of questions. One is completely unrelated knowledge, for which we use the prompt “The capital city of America is”. The other type has the same relation as the edited knowledge but a different subject. For example, if the edited knowledge is (John, father, Bob), a specificity question could be (Paul, father, Eugène).
@@ Appendix C More Results with Varying Batch Sizes
It shows clearly that MEMIT-based methods suffers from same subject issue, while methods like FT do not.